# Effective models for dense vortex lattices in the Kitaev honeycomb model

### 22 Dec 2023 v1

@@ IV.2 Dual even case
In this section we introduce a low-energy model for dual even vortex configurations. The size of the HUC is double that of the GUC, and we set $\bm{a}_{1}=2\bm{A}_{1}$, $\bm{a}_{2}=\bm{A}_{2}$, where the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ now generate the dual vortex lattice. Each of these isolated $w_{p}=+1$ plaquettes (i.e. a “dual vortex”) hosts two Majorana modes at finite energy ±$\pm\epsilon$. The splitting $2\epsilon$ has been calculated numerically and plotted in Fig. 6 of Ref. [15]. It vanishes in the $J=0$ limit in which the two triangular sublattices of the honeycomb lattice are decoupled (see the discussion on the vortex-full sector in Sec. II). This means that the two Majorana modes in the same “dual vortex” are uncoupled and therefore at zero energy. Inspired by this $J=0$ limit and by the effective model in the direct case [see Eq. (3)], we construct a tight-binding model that features two “layers” (or flavors) of Majorana modes. They tunnel between the dual vortices according to the following effective Hamiltonian:
@@
Figure 6: Dual even: Unit cell of the effective model introduced in Eq. (16). Each “dual vortex” hosts two Majorana modes, which we interpret as two “layers” that we color white and light blue. Here, $\epsilon$ represents the on-site coupling of two modes on the same site, and the blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest and next-nearest neighbor hopping terms, respectively. The pattern of these gauge variables for both layers is the same as that given in Fig. 4.
Figure 7: Dual even: (a) Phase diagram and Chern numbers $\nu_{\rm v}$ of the effective model given in Eq. (16), where $t_{1}<0$. $E_{\rm min}(t_{1},t_{2})$ and $E_{\rm max}(t_{1},t_{2})$ are defined in Eq. (14). (b) Comparison of the IDOS of the effective model and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here, $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.122$ and the parameters are $\epsilon=0.370$, $t_{1}=-0.052$, and $t_{2}=0.004$. These parameter values are indicated by the red dot in the phase diagram of (a), implying that $\nu_{\rm v}=0$, in agreement with the microscopic model.
In this construction, each isolated “dual vortex” in Fig. 1(b) hosts two Majorana modes: one belonging to $\alpha_{j}$, and one belonging to $\beta_{j}$. The on-site coupling between them is given by $\epsilon$, and at the current level we do not allow any inter-layer tunneling between neighboring dual-vortex sites; nearest and next-nearest neighbor tunneling occurs only between sites of the same layer. Including such terms would violate the $\mathcal{P}^{2}=+\mathbbm{1}$ symmetry of the microscopic model. As before, $s_{jk}^{(\alpha)}=\pm 1$ and $s_{jk}^{(\beta)}=\pm 1$ are $\mathbb{Z}_{2}$ gauge variables that determine the total enclosed flux of triangular circuits of dual vortices. From the $P_{w}=-1$ symmetry [which can be seen in Fig. 1(b)], the total flux enclosed when hopping counter-clockwise on a nearest neighbor circuit of $t_{1}$ links is $\phi_{1}=-\pi/2$, the flux enclosed in a next-nearest neighbor circuit of $t_{2}$ links is $\phi_{2}=\pi/2$, and the flux enclosed in a circuit composed of both $t_{1}$ and $t_{2}$ links is $\phi_{1,2}=-\pi/2$. We may observe that this is simply the opposite of the flux pattern of the previous section. Noting that exchanging $\kappa\to-\kappa$ reverses the flux, we may adopt the same pattern of $s_{jk}$ variables defined in Fig. 4(b) for both layers of Eq. (15). That is, we set $s_{jk}^{(\alpha)}=s_{jk}^{(\beta)}$. Doing so allows us to Fourier transform the Hamiltonian of Eq. (15) to find
@@
In Fig. 7(b) we plot the IDOS of the low-energy bands of $H$ along with that of the effective Hamiltonian $H_{2}$. As before, we fit the parameters $\epsilon$, $t_{1}$, and $t_{2}$ by matching the IDOS and Chern number with Eq. (1).
@@ IV.3 Dual odd case
We wish to construct a tight-binding model that features two layers of Majorana modes that tunnel between dual vortices located. However, in anticipation of the $\mathcal{P}^{2}=-\mathbbm{1}$ symmetry constraint, this time we shall allow inter-layer tunneling between neighboring dual-vortex sites. The Hamiltonian is then
@@
Here, $(t_{1},t_{2})\in\mathbb{R}^{2}$ represent the hopping amplitudes of modes between nearest and next-nearest neighbor sites in the same layer, respectively. In contrast, the new terms $(t_{\perp 1},t_{\perp 2})\in\mathbb{R}^{2}$ represent the inter-layer hopping amplitudes between nearest and next-nearest neighbors, respectively. This Hamiltonian then has the four sets of gauge variables: $s^{(\alpha)}_{jk}$, $s^{(\beta)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$. To obey the $\mathcal{P}^{2}=-\mathbbm{1}$ symmetry we must set $s^{(\beta)}_{jk}=-s^{(\alpha)}_{jk}$, which leaves us with three available sets of gauge variables to choose in order to analyze the model, along with the signs of the hopping parameters and $\epsilon$.
Figure 8: Dual odd: Unit cell of the effective model defined in Eq. (19). (a) The blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest $(t_{1})$ and next-nearest neighbor $(t_{2})$ intra-layer hopping terms, respectively. The upper layer corresponds to the gauge variables $s^{(\alpha)}_{jk}$, while the lower layer corresponds to the gauge variables $s^{(\beta)}_{jk}=-s^{(\alpha)}_{jk}$. (b) Schematic of the inter-layer hopping amplitudes in Eq. (19) (top view). The terms $t_{\perp 1}$ and $t_{\perp 2}$ are colored blue and red, respectively. Solid arrows refer to the choice of signs for the $\widetilde{s}^{(\perp)}_{jk}$ gauge variables, while dashed arrows refer to $s^{(\perp)}_{jk}$.
Figure 9: Dual odd: Phase diagram and Chern numbers $\nu_{\rm v}$ of the model given in Eq. (19). (a) $t_{1}=0$ and $t_{2}=t_{\perp 2}$. The phase boundaries are $t_{\perp 1}=\epsilon/6,\epsilon,-\epsilon/2$, and $-\epsilon/3$. (b) $t_{2}=0$ and $t_{1}=t_{\perp 2}$. The phase boundaries are $t_{\perp 1}=\epsilon/6$ and ±$\pm\epsilon/2$.
The original honeycomb model in Eq. (1) shows that for this family the accumulated flux in a triangular circuit of nearest-neighbor dual vortices is found to be always either $\phi=0$ or $\phi=\pi$. Unlike in the previous two sections, the gauge variables of Eq. (18) cannot be chosen in such a way that recreates this pattern. Indeed, for Majorana fermions, triangular circuits always carry a flux of ±$\pm\pi/2$. However, as schematically shown in Fig. 8, the size of the HUC implies that all gauge choices of the $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$ variables result in patterns of alternating ±$\pm\pi/2$ flux within the nearest neighbor triangular circuits, which is the closest a model of Majorana fermions can be to obtaining a flux of 00 or $\pi$. As such, we are therefore unrestricted in our choice of gauge variables. Taking the gauge choices for $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$ defined in Fig. 8, we may Fourier transform the Hamiltonian to find
@@
The nearest neighbor directions are defined as $\bm{a}_{1}$, $\bm{a}_{2}$, and $\bm{a}_{3}=\bm{a}_{2}-\bm{a}_{1}$, while the next-nearest neighbor directions are defined as $\bm{b}_{1}=\bm{a}_{1}+\bm{a}_{2}$, $\bm{b}_{2}=\bm{a}_{2}+\bm{a}_{3}$, and $\bm{b}_{3}=\bm{a}_{3}-\bm{a}_{1}$. The above Majorana Hamiltonian in Eq. (26) can be transformed into a Bogoliubov-de Gennes Hamiltonian by a $\bm{k}$-independent unitary transformation, showing that the model is close to a $p_{x}+ip_{y}$ superconductor on a triangular lattice. This could have been anticipated given that a “dual vortex” hosts a single complex fermionic mode, that the energy spectrum must have particle-hole symmetry, and that the gapped bands carry nonzero Chern numbers.
@@
Figure 10: Dual odd: Comparison of the IDOS of the effective model given in Eq. (19) and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here, $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.038$ and the parameters are $\epsilon=0.395$, $t_{1}=0.095$, $t_{\perp 1}=-0.043$, and all the other terms zero (note that these values do not correspond to one of the phase diagrams given in Fig. 9). For an effective model given by these parameters we calculate the vortex Chern number to be $\nu_{\rm v}=0$, in agreement with the microscopic model.
@@
In Figs. 9(a) and (b) we plot the phase diagrams of this model as $t_{1}=0$, $t_{2}=t_{\perp 2}$ and $t_{2}=0$, $t_{1}=t_{\perp 2}$, respectively, which collectively exhibit the Chern numbers $\nu_{\rm v}=0,\pm 1,\pm 2$, and ±$\pm 3$. Thus, knowing that the vortex-full background exhibits $\nu_{\rm b}=\pm 2$, the total Chern numbers arising from the above effective model are $\nu=0,\pm 1,\pm 2,\pm 3,\pm 4$ and ±$\pm 5$. This is in agreement with what is observed in Ref. [15], which, for this family, showcases all of these Chern numbers except $\nu=\pm 4$. We also note that the above model is just one of several possible choices for the gauge variables $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$, and other choices may give rise to additional Chern numbers. In total, the model in Eq. (26) is determined by the five parameters $\epsilon,t_{1},t_{2},t_{\perp 1}$, and $t_{\perp 2}$, along with the choice of four signs. These are the signs in front of $\cos(\bm{k}\cdot\bm{a}_{2})$, $\cos(\bm{k}\cdot\bm{a}_{3})$, $\sin(\bm{k}\cdot\bm{b}_{2})$ and $\sin(\bm{k}\cdot\bm{b}_{3})$ in Eq. (26). All other signs can be absorbed into the five parameters.
@@
In Fig. 10 we plot the IDOS of the low-energy bands of the Kitaev honeycomb model in Eq. (1) along with that of the effective model in Eq. (19). The five parameters $\epsilon$, $t_{1}$, $t_{2}$, $t_{\perp 1}$, and $t_{\perp 2}$ are found by assuming a relevant gauge and matching the IDOS and Chern number with Eq. (1). If the Chern number of the model is not known beforehand, then multiple sets of these parameters may give rise to the same IDOS. In this case, unlike the previous two symmetry families, there is no clear way to predict the relative magnitudes of these five parameters by studying a two “dual vortex” problem numerically.
@@ V Discussion and perspectives
In this article, we extended the results of previous works by introducing low-energy models that can describe both sparse triangular (direct) and dual vortex configurations in Kitaev’s honeycomb model. In the direct case, featuring a triangular pattern of vortices against a vortex-free background, the behavior of the Kitaev honeycomb model may be interpreted as arising from a combination of the background and of the tunneling of single MZMs between vortex sites. In the dual case, a triangular pattern of vortex-absences is instead superimposed on top of a vortex-full background. These isolated “dual vortices” each contain a pair of coupled Majorana modes, and we have shown that the properties of the Kitaev honeycomb model may be interpreted as arising from a combination of the vortex-full background and of two “layers” of finite-energy Majorana modes that tunnel to other dual-vortex sites in either an intra-layer or inter-layer fashion.
In both the direct and dual cases the phase diagram of the Kitaev honeycomb model is controlled by the TRS breaking term $\kappa$ that we allow to be arbitrarily strong compared to the nearest neighbor interaction $J$. For certain values of $\kappa/J$, the low-lying energy bands of the microscopic model are well-separated from the high-energy bands, and in these cases we have demonstrated that the low-energy bands may be accurately modeled with our effective descriptions. In addition, the high-energy bands in the microscopic model resemble the spectra of the vortex-free or vortex-full backgrounds. Furthermore, for both the direct and dual configurations, we have found that the total Chern number of the microscopic model $\nu$ is equal to the sum of the background bands’ Chern number $\nu_{\rm b}$ and the low-energy bands’ Chern number $\nu_{\rm v}$, i.e., $\nu=\nu_{\rm b}+\nu_{\rm v}$. The Chern numbers predicted by our effective models are in very good agreement with what has been observed in Ref. [15].
@@
In addition, there are several open questions raised by this work that should be addressed. For example, is it possible to find a phase with Chern number $\nu=\pm 7$ mod 16 by considering nontriangular vortex configurations [15]? Are there models that realize the fourth family (see Table 2) of inversion-symmetric and periodic lattices that is expected to feature extended gapless phases and even or odd Chern numbers, or is there a general argument showing that this is impossible? What happens in cases where inversion symmetry does not commute with the translation operators defining the HUC? More generally, can we use other crystalline symmetries than inversion to classify all type of behavior (parity of Chern number, nature of gapless phases) in the Kitaev model in different vortex sectors and on different lattices (the honeycomb but also the square-octagon [23] or the triangle-dodecagon lattice [26, 27, 24, 28], for example)?

### 05 Mar 2024 v2

@@ IV.2 Dual even case
In this section we introduce a low-energy model for dual even vortex configurations. The size of the HUC is double that of the GUC, and we set $\bm{a}_{1}=2\bm{A}_{1}$, $\bm{a}_{2}=\bm{A}_{2}$, where the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ now generate the dual vortex lattice. Each of these isolated $w_{p}=+1$ plaquettes (i.e., a “dual vortex”) hosts two Majorana modes at finite energy ±$\pm\epsilon$. The splitting $2\epsilon$ has been calculated numerically and plotted in Fig. 6 of Ref. [17]. It vanishes in the $J=0$ limit in which the two triangular sublattices of the honeycomb lattice are decoupled (see the discussion on the vortex-full sector in Sec. II). This means that the two Majorana modes in the same “dual vortex” are uncoupled and therefore at zero energy. Inspired by this $J=0$ limit and by the effective model in the direct case [see Eq. (3)], we construct a tight-binding model that features two “layers” (or flavors) of Majorana modes. They tunnel between the dual vortices according to the following effective Hamiltonian:
@@
Figure 6: Dual even. Unit cell of the effective model introduced in Eq. (16). Each “dual vortex” hosts two Majorana modes, which we interpret as two “layers” that we color white and light blue. Here, $\epsilon$ represents the on-site coupling of two modes on the same site, and the blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest and next-nearest-neighbor hopping terms, respectively. The pattern of these gauge variables for both layers is the same as that given in Fig. 4.
Figure 7: Dual even. (a) Phase diagram and Chern numbers $\nu_{\rm v}$ of the effective model given in Eq. (16), where $t_{1}<0$. $E_{\rm min}(t_{1},t_{2})$ and $E_{\rm max}(t_{1},t_{2})$ are defined in Eq. (14). (b) Comparison of the IDOS of the effective model and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here, $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.122$ and the parameters are $\epsilon=0.370$, $t_{1}=-0.052$, and $t_{2}=0.004$. These parameter values are indicated by the red dot in the phase diagram of (a), implying that $\nu_{\rm v}=0$, in agreement with the microscopic model.
In this construction, each isolated “dual vortex” in Fig. 1(b) hosts two Majorana modes: one belonging to $\alpha_{j}$, and one belonging to $\beta_{j}$. The on-site coupling between them is given by $\epsilon$, and at the current level we do not allow any interlayer tunneling between neighboring dual-vortex sites; nearest and next-nearest-neighbor tunneling occurs only between sites of the same layer. Including such terms would violate the $\mathcal{P}^{2}=+\mathbbm{1}$ symmetry of the microscopic model. As before, $s_{jk}^{(\alpha)}=\pm 1$ and $s_{jk}^{(\beta)}=\pm 1$ are $\mathbb{Z}_{2}$ gauge variables that determine the total enclosed flux of triangular circuits of dual vortices. From the $P_{w}=-1$ symmetry [which can be seen in Fig. 1(b)], the total flux enclosed when hopping counter-clockwise on a nearest-neighbor circuit of $t_{1}$ links is $\phi_{1}=-\pi/2$, the flux enclosed in a next-nearest-neighbor circuit of $t_{2}$ links is $\phi_{2}=\pi/2$, and the flux enclosed in a circuit composed of both $t_{1}$ and $t_{2}$ links is $\phi_{1,2}=-\pi/2$. We may observe that this is simply the opposite of the flux pattern of the previous section. Noting that exchanging $\kappa\to-\kappa$ reverses the flux, we may adopt the same pattern of $s_{jk}$ variables defined in Fig. 4(b) for both layers of Eq. (15). That is, we set $s_{jk}^{(\alpha)}=s_{jk}^{(\beta)}$. Doing so allows us to Fourier transform the Hamiltonian of Eq. (15) to find
@@
In Fig. 7(b), we plot the IDOS of the low-energy bands of $H$ along with that of the effective Hamiltonian $H_{2}$. As before, we fit the parameters $\epsilon$, $t_{1}$, and $t_{2}$ by matching the IDOS and Chern number with Eq. (1).
@@ IV.3 Dual odd case
We wish to construct a tight-binding model that features two layers of Majorana modes that tunnel between dual vortices located. However, in anticipation of the $\mathcal{P}^{2}=-\mathbbm{1}$ symmetry constraint, this time we shall allow interlayer tunneling between neighboring dual-vortex sites. The Hamiltonian is then
@@
Here, $(t_{1},t_{2})\in\mathbb{R}^{2}$ represent the hopping amplitudes of modes between nearest and next-nearest-neighbor sites in the same layer, respectively. In contrast, the new terms $(t_{\perp 1},t_{\perp 2})\in\mathbb{R}^{2}$ represent the interlayer hopping amplitudes between nearest and next-nearest neighbors, respectively. This Hamiltonian then has the four sets of gauge variables: $s^{(\alpha)}_{jk}$, $s^{(\beta)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$. To obey the $\mathcal{P}^{2}=-\mathbbm{1}$ symmetry we must set $s^{(\beta)}_{jk}=-s^{(\alpha)}_{jk}$, which leaves us with three available sets of gauge variables to choose in order to analyze the model, along with the signs of the hopping parameters and $\epsilon$.
Figure 8: Dual odd. Unit cell of the effective model defined in Eq. (19). (a) The blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest $(t_{1})$ and next-nearest-neighbor $(t_{2})$ intralayer hopping terms, respectively. The upper layer corresponds to the gauge variables $s^{(\alpha)}_{jk}$, while the lower layer corresponds to the gauge variables $s^{(\beta)}_{jk}=-s^{(\alpha)}_{jk}$. (b) Schematic of the interlayer hopping amplitudes in Eq. (19) (top view). The terms $t_{\perp 1}$ and $t_{\perp 2}$ are colored blue and red, respectively. Solid arrows refer to the choice of signs for the $\widetilde{s}^{(\perp)}_{jk}$ gauge variables, while dashed arrows refer to $s^{(\perp)}_{jk}$.
Figure 9: Dual odd. Phase diagram and Chern numbers $\nu_{\rm v}$ of the model given in Eq. (19). (a) $t_{1}=0$ and $t_{2}=t_{\perp 2}$. The phase boundaries are $t_{\perp 1}=\epsilon/6,\epsilon,-\epsilon/2$, and $-\epsilon/3$. (b) $t_{2}=0$ and $t_{1}=t_{\perp 2}$. The phase boundaries are $t_{\perp 1}=\epsilon/6$ and ±$\pm\epsilon/2$.
The original honeycomb model in Eq. (1) shows that for this family the accumulated flux in a triangular circuit of nearest-neighbor dual vortices is found to be always either $\phi=0$ or $\phi=\pi$. Unlike in the previous two sections, the gauge variables of Eq. (18) cannot be chosen in such a way that recreates this pattern. Indeed, for Majorana fermions, triangular circuits always carry a flux of ±$\pm\pi/2$. However, as schematically shown in Fig. 8, the size of the HUC implies that all gauge choices of the $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$ variables result in patterns of alternating ±$\pm\pi/2$ flux within the nearest-neighbor triangular circuits, which is the closest a model of Majorana fermions can be to obtaining a flux of 00 or $\pi$. As such, we are therefore unrestricted in our choice of gauge variables. Taking the gauge choices for $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$ defined in Fig. 8, we may Fourier transform the Hamiltonian to find
@@
The nearest-neighbor directions are defined as $\bm{a}_{1}$, $\bm{a}_{2}$, and $\bm{a}_{3}=\bm{a}_{2}-\bm{a}_{1}$, while the next-nearest-neighbor directions are defined as $\bm{b}_{1}=\bm{a}_{1}+\bm{a}_{2}$, $\bm{b}_{2}=\bm{a}_{2}+\bm{a}_{3}$, and $\bm{b}_{3}=\bm{a}_{3}-\bm{a}_{1}$. The above Majorana Hamiltonian in Eq. (26) can be transformed into a Bogoliubov-de Gennes Hamiltonian by a $\bm{k}$-independent unitary transformation, showing that the model is close to a $p_{x}+ip_{y}$ superconductor on a triangular lattice. This could have been anticipated given that a “dual vortex” hosts a single complex fermionic mode, that the energy spectrum must have particle-hole symmetry, and that the gapped bands carry nonzero Chern numbers.
@@
Figure 10: Dual odd. Comparison of the IDOS of the effective model given in Eq. (19) and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here, $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.038$ and the parameters are $\epsilon=0.395$, $t_{1}=0.095$, $t_{\perp 1}=-0.043$, and all the other terms zero (note that these values do not correspond to one of the phase diagrams given in Fig. 9). For an effective model given by these parameters we calculate the vortex Chern number to be $\nu_{\rm v}=0$, in agreement with the microscopic model.
@@
In Figs. 9(a) and 9(b), we plot the phase diagrams of this model as $t_{1}=0$, $t_{2}=t_{\perp 2}$ and $t_{2}=0$, $t_{1}=t_{\perp 2}$, respectively, which collectively exhibit the Chern numbers $\nu_{\rm v}=0,\pm 1,\pm 2$, and ±$\pm 3$. Thus, knowing that the vortex-full background exhibits $\nu_{\rm b}=\pm 2$, the total Chern numbers arising from the above effective model are $\nu=0,\pm 1,\pm 2,\pm 3,\pm 4$ and ±$\pm 5$. This is in agreement with what is observed in Ref. [17], which, for this family, showcases all of these Chern numbers except $\nu=\pm 4$. We also note that the above model is just one of several possible choices for the gauge variables $s^{(\alpha)}_{jk}$, $s^{(\perp)}_{jk}$, and $\widetilde{s}^{(\perp)}_{jk}$, and other choices may give rise to additional Chern numbers. In total, the model in Eq. (26) is determined by the five parameters $\epsilon,t_{1},t_{2},t_{\perp 1}$, and $t_{\perp 2}$, along with the choice of four signs. These are the signs in front of $\cos(\bm{k}\cdot\bm{a}_{2})$, $\cos(\bm{k}\cdot\bm{a}_{3})$, $\sin(\bm{k}\cdot\bm{b}_{2})$ and $\sin(\bm{k}\cdot\bm{b}_{3})$ in Eq. (26). All other signs can be absorbed into the five parameters.
@@
In Fig. 10, we plot the IDOS of the low-energy bands of the Kitaev honeycomb model in Eq. (1) along with that of the effective model in Eq. (19). The five parameters $\epsilon$, $t_{1}$, $t_{2}$, $t_{\perp 1}$, and $t_{\perp 2}$ are found by assuming a relevant gauge and matching the IDOS and Chern number with Eq. (1). If the Chern number of the model is not known beforehand, then multiple sets of these parameters may give rise to the same IDOS. In this case, unlike the previous two symmetry families, there is no clear way to predict the relative magnitudes of these five parameters by studying a two “dual vortex” problem numerically.
@@ V Discussion and perspectives
In this paper, we extended the results of previous works by introducing low-energy models that can describe both sparse triangular (direct) and dual vortex configurations in Kitaev’s honeycomb model. In the direct case, featuring a triangular pattern of vortices against a vortex-free background, the behavior of the Kitaev honeycomb model may be interpreted as arising from a combination of the background and of the tunneling of single MZMs between vortex sites. In the dual case, a triangular pattern of vortex-absences is instead superimposed on top of a vortex-full background. These isolated “dual vortices” each contain a pair of coupled Majorana modes, and we have shown that the properties of the Kitaev honeycomb model may be interpreted as arising from a combination of the vortex-full background and of two “layers” of finite-energy Majorana modes that tunnel to other dual-vortex sites in either an intralayer or interlayer fashion.
In both the direct and dual cases, the phase diagram of the Kitaev honeycomb model is controlled by the TRS breaking term $\kappa$ that we allow to be arbitrarily strong compared to the nearest-neighbor interaction $J$. For certain values of $\kappa/J$, the low-lying energy bands of the microscopic model are well-separated from the high-energy bands, and in these cases we have demonstrated that the low-energy bands may be accurately modeled with our effective descriptions. In addition, the high-energy bands in the microscopic model resemble the spectra of the vortex-free or vortex-full backgrounds. Furthermore, for both the direct and dual configurations, we have found that the total Chern number of the microscopic model $\nu$ is equal to the sum of the background bands’ Chern number $\nu_{\rm b}$ and the low-energy bands’ Chern number $\nu_{\rm v}$, i.e., $\nu=\nu_{\rm b}+\nu_{\rm v}$. The Chern numbers predicted by our effective models are in very good agreement with what has been observed in Ref. [17].
@@
In addition, there are several open questions raised by this work that should be addressed. For example, is it possible to find a phase with Chern number $\nu=\pm 7$ mod 16 by considering nontriangular vortex configurations [17]? Are there models that realize the fourth family (see Table 2) of inversion-symmetric and periodic lattices that is expected to feature extended gapless phases and even or odd Chern numbers, or is there a general argument showing that this is impossible? What happens in cases where inversion symmetry does not commute with the translation operators defining the HUC? More generally, can we use other crystalline symmetries than inversion to classify all type of behavior (parity of Chern number, nature of gapless phases) in the Kitaev model in different vortex sectors and on different lattices (the honeycomb but also the square-octagon [26] or the triangle-dodecagon [29, 30, 27, 31] or the pentaheptite lattice [18], for example)?